OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the points (−3, 2) and (2, 1). y + 3 = −5(x − 2) y − 2 = −5(x + 3) y + 3 = −1/5(x − 2) y − 2 = −1/5(x + 3)

OpenStudy (lgbasallote):

use the formula \[\large y - y_1 = \frac{y_2 - y_1}{x_2 - x_1} (x - x_1)\]

OpenStudy (anonymous):

line direction vector (2-(-3),1-2)=(5,-1) so slope =-1/5 so it's y + 3 = −1/5(x − 2)